OpenStudy (anonymous):

An auto manufacturer sends cars from two plants, I and II, to dealerships A and B, located in a mid-western city. Plant I has a total of 28 cars to send, and plant II has 8. Dealer A needs 20 cars, and dealer B needs 16. Transportation costs based on the distance of each dealership from each plant are $220 from I to A, $300 from I to B, $400 from II to A, and $180 from II to B. The manufacturer wants to limit transportation costs to $10,640. How many cars should be sent from each plant to each of the two dealerships? (Hints: Define each unknown value with a variable. There should be four variables in all

OpenStudy (amistre64):

Plant I has a total of 28 cars Plant II has 8 A needs 20 cars B needs 16. **************** Transportation costs (is this per car or just for the trip?) 220 from I to A, 300 from I to B, 400 from II to A, 180 from II to B. The manufacturer wants to limit transportation costs to $10,640.

OpenStudy (anonymous):

yes.can you help @amistre64

OpenStudy (amistre64):

can you tell me what math this is for? linear algebra and matrixes by chance? or something else

OpenStudy (anonymous):

linear equation

OpenStudy (anonymous):

systems of equation

OpenStudy (amistre64):

im thinking ...... if we divvy up the cars to the dealers we get A = 20 = mi + nii B = 16 = (28-m)i + (8-n)ii if we calculate the cost per car A = 220m + 400n B = 300(28-m) + 180(8-n) such that 220m + 400n + 300(28-m) + 180(8-n) <= 10640

OpenStudy (amistre64):

220m + 400n + 300(28) -300m + 180(8) -180n <= 10640 320n -80m <= 10640 - 180(8) - 300(28) 4n - m <= 10 4n - 10 <= m 20 = (4n-10) + n 30 = 5n n = 6 A = 20 = 14i + 6ii B = 16 = 14i + 2ii but this is just some rough thinking and prolly should be modified

OpenStudy (anonymous):

i thought the question said there need to be four variables in all

OpenStudy (amistre64):

thats what the hint said ... if youve got another way to figure this out, then by all means .....

OpenStudy (anonymous):

please what is the answer now.is the answer 6?

OpenStudy (amistre64):

lets see how this applies to price 220(14) from I to A, 300(6) from I to B, 400(14) from II to A, 180(2) from II to B. gives me 10840 so its a little high, but close

OpenStudy (amistre64):

i dont know what the answer is ... but i have an idea on how to go about it now

OpenStudy (anonymous):

ok then please do help :-) waiting

OpenStudy (amistre64):

to reduce the cost; lets adjust some values 220(15) from I to A, 300(6) from I to B, 400(13) from II to A, 180(2) from II to B. gets me to 10660 still a wee bit too high for 10640 --------------------------------- 220(15) from I to A, 300(5) from I to B, 400(13) from II to A, 180(3) from II to B. gets me to 10540 which fulfills the costs portion 10640 -------------------------------------- so the question i have is .... does the cost have to be a minimum cost? or just less than or equal to 10640?

OpenStudy (anonymous):

just less than or equal to i guez

OpenStudy (amistre64):

220(20) from I to A, 300(0) from I to B, 400(8) from II to A, 180(8) from II to B. gets me to 9040

OpenStudy (amistre64):

im mixing my data up ... so thats not correct either

OpenStudy (anonymous):

lol.ok.i want to know how u getting the numbers in the bracket?

OpenStudy (amistre64):

we have 28 cars from i to send; it costs 220(A) + 300(B) to divvy up 28 cars between A and B we have 8 cars from ii to send; it costs 400(A) + 180(B) to divvy up 8 cars between A and B

OpenStudy (amistre64):

A wants 20 total B wants 16 total if we go with just the cheapest stuff first: A B 28: 220(20) + 300(8) 8: 400(0) + 180(8) ------------------- 20 16

OpenStudy (amistre64):

thats a total charge of 8240

OpenStudy (anonymous):

yep.so how many cars should be sent from each plant of the two dealerships

OpenStudy (amistre64):

thats what the nifty setup tells us .... A B from i 28: 220(20i) + 300(8i) from ii 8: 400(0ii) + 180(8ii) ------------------- 20 16

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

are you gud in sloving a 3x3 larger system of equation

OpenStudy (amistre64):

theres prolly a way to do it with 4 variables and whatnot ... but i just cant see it that way :/ 3x3 is best solved with a matrix

OpenStudy (anonymous):

should i send it over to you and then see to it for me?

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

you can post it right here and i will be able to view it ... i got no idea of i can solve it yet or not :)

OpenStudy (anonymous):

dollars on his Mastercard, y dollars on his Visa, and z dollars on his American Express card. These amounts satisfy the following equations. 1.18x+1.15y+1.09z=11,244.25 3.54x-.55y+.27z=3,732.75 .12x+.10y+.06z=829.50 How much did the owner borrow on each card?

OpenStudy (anonymous):

lol.i hope u can @amistre64

OpenStudy (amistre64):

piece of cake :)

OpenStudy (anonymous):

hahaha

OpenStudy (amistre64):

rref{{1.18,1.15,1.09,11244.25},{3.54,-.55,.27,3732.75},{.12,.10,.06,829.50}} http://www.wolframalpha.com/input/?i=rref%7B%7B1.18%2C1.15%2C1.09%2C11244.25%7D%2C%7B3.54%2C-.55%2C.27%2C3732.75%7D%2C%7B.12%2C.10%2C.06%2C829.50%7D%7D

OpenStudy (amistre64):

the last column top to bottom is the values for xyz

OpenStudy (amistre64):

if your intent is to do it by hand .... then no, i simply dont have time to do it :)